1. A satellite is orbiting Earth on a path described by x2+y2=45 000 000. A second satellite is in the same plane and is occupying the point (12 504, 16 050). Determine if the orbit is narrower, the same, or wider than the first satellite's.
2. Write the equation for a circle centered at the origin with x-intercepts of (-9,0) and (9,0).Use ^2 for squared, ^3 for cubed, etc.

Answer by ikleyn(52781)

.
1. A satellite is orbiting Earth on a path described by x2+y2=45 000 000.
A second satellite is in the same plane and is occupying the point (12 504, 16 050).
Determine if the orbit is narrower, the same, or wider than the first satellite's.
2. Write the equation for a circle centered at the origin with x-intercepts of (-9,0) and (9,0).
Use ^2 for squared, ^3 for cubed, etc.
~~~~~~~~~~~~~~~~~~~~~~~~

Regarding first question, it is posed incorrectly.


It is posed incorrectly, because the orbit of the second satellite is not defined.
It can be an elliptical curve, which is partly inside of the first orbit and
partly outside it.


The correct question would be "whether the given point lies inside or outside of the given circle ?".

To answer it, calculate the square of the distance of the given point from the origin 
and compare it with 45,000,000.



Regarding the second question, the sough equation is

    x^2 + y^2 = 9^2.

Solved.

---------------

The rules of this forum PROHIBIT posting more than one question per post.


The rules are at this web-page

https://www.algebra.com/tutors/students/ask.mpl?action=ask_question&topic=Equations&return_url=http://www.algebra.com/algebra/homework/equations/

from which you post your problems.


Please respect the rules.